Find the value of X and Y in each of the following
3x(2+5i)-4(2x-3yi)=12+13i

Answer by Alan3354(21584)

3x(2+5i)-4(2x-3yi)=12+13i
6x + 15ix - 8x + 12iy = 12 + 13i
-2x + 15ix + 12iy = 12 + 13i
---------
-2x = 12
x = -6
-----------
i(15x+12y) = 13i
15x+12y = 13
-90+12y = 13
12y = 103
y = 103/12